White Company has two departments, Cutting and Finishing. The company uses a job-order costing system and computes a predetermined overhead rate in each department. The Cutting Department bases its rate on machine-hours, and the Finishing Department bases its rate on direct labor-hours. At the beginning of the year, the company made the following estimates:

Department
Cutting Finishing
Direct labor-hours 8,100 66,000
Machine-hours 57,300 1,400
Total fixed manufacturing overhead cost $ 400,000 $ 478,000
Variable manufacturing overhead per machine-hour $ 3.00
Variable manufacturing overhead per direct labor-hour $ 2.75

Required:

1. Compute the predetermined overhead rate for each department.

2. The job cost sheet for Job 203, which was started and completed during the year, showed the following:

Department
Cutting Finishing
Direct labor-hours 5 11
Machine-hours 90 6
Direct materials $ 790 $ 360
Direct labor cost $ 95 $ 209

Using the predetermined overhead rates that you computed in requirement (1), compute the total manufacturing cost assigned to Job 203.

3. Would you expect substantially different amounts of overhead cost to be assigned to some jobs if the company used a plantwide predetermined overhead rate based on direct labor-hours, rather than using departmental rates?
Answer #1

1) Departmental overhead rate

Cutting overhead rate = 400000/57300+3 = 9.98 per MH

Finishing overhead rate = 478000/66000+2.75 = 9.99 per DLH

2) Total Manufacturing cost

Cutting Finishing Total
Direct material 790 360 1150
Direct labor 95 209 304
Overhead 898.20 109.89 1008.09
Total 1783.20 678.89 2462.09 or 2462

3) Yes